Use trigonometric identities to solve sin(theta)+(sqrt3)sin(theta)=0 exactly for 0 is less than or equal to theta which is less than 2pi;. If there is more than one answer, enter your answers as a comma separated list.
Thank you

Found 2 solutions by lwsshak3, ikleyn:
Answer by lwsshak3(11628)

Use trigonometric identities to solve sin(theta)+(sqrt3)sin(theta)=0 exactly for 0 is less than or equal to theta which is less than 2pi;. If there is more than one answer, enter your answers as a comma separated list.
***
sinx+√3sinx=0
sinx≠-√3sinx
no solution


Answer by ikleyn(53875)

.
Use trigonometric identities to solve sin(theta)+(sqrt3)sin(theta)=0 exactly for 0 is less than or equal to theta which is less than 2pi;.
If there is more than one answer, enter your answers as a comma separated list.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution and the answer in the post by @lwsshak3 both are incorrect.
        Find my correct solution below.


sin(x) + √3*sin(x) = 0
sin(x)*(1 + √3) = 0.

It implies

sin(x) = 0,
x = 0 in the given interval.